OpenStudy (anonymous):

The graph below shows three functions: Graph of function f of x equals 0.5 to the power x. Graph of function g of x is a line segment with holes at endpoints connecting negative 1.8, negative 3 and 1, 3.8. Graph of p of x is a line segment connecting negative 1.7, 2.3 and 1, negative 4.2. Which function has all real numbers as its domain? Only f(x) Both g(x) and p(x) Both f(x) and g(x) Only p(x)

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (cbarredo1):

You know what are real numbers, right?

OpenStudy (anonymous):

yes @CBARREDO1

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (cbarredo1):

Ok, I'm going to give you the formulas \[f(x)=0.5^x\] \[g(x)= (-1.8,-3) (1,3.8)\] \[p(x)=(1.7,2.3) (1,-4.2)\]

OpenStudy (cbarredo1):

Of this formulas which one contains only real numbers?

OpenStudy (anonymous):

1

OpenStudy (cbarredo1):

There xD

OpenStudy (anonymous):

@CBARREDO1 it is A?

OpenStudy (cbarredo1):

I think so :)

OpenStudy (anonymous):

1 more please

OpenStudy (anonymous):

Which of the following shows the correct steps to find the value of 16 to the power of 1 over 4?

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (cbarredo1):

What are the options?

OpenStudy (anonymous):

16 to the power of 1 over 4 equals 2 to the power of 4 to the power of 1 over 4 equals 2 to the power of 4 multiplied by 1 over 4 equals 2 16 to the power of 1 over 4 equals 4 to the power of 4 to the power of 1 over 4 equals 4 to the power of 4 multiplied by 1 over 4 equals 4 16 to the power of 1 over 4 equals 2 to the power of 8 to the power of 1 over 4 equals 8 to the power of 8 multiplied by 1 over 4 equals 4 16 to the power of 1 over 4 equals 8 to the power of 2 to the power of 1 over 4 equals 2 to the power of 2 multiplied by 1 over 4 equals 8

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (cbarredo1):

Which of the following shows the correct steps to find the value of 16 to the power of 1 over 4? \[(16)^1/4=(2^3)^1/4=2^3*1/4=2^1=2\]

OpenStudy (cbarredo1):

power of 2 sorry

OpenStudy (cbarredo1):

D is the answer

OpenStudy (anonymous):

The amount of a radioactive material changes with time. The table below shows the amount of radioactive material f(t) left after time t: t(hours) 0 1 2 f(t) 180 90 45 Which exponential function best represents the relationship between f(t) and t? f(t) = 0.5(180)t f(t) = 180(0.25)t f(t) = 180(0.5)t f(t) = 0.5(50)t

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (cbarredo1):

Who would think I would end up doing algebra in the summer? I don't get paid enough for this xD

OpenStudy (anonymous):

XD

OpenStudy (cbarredo1):

f(t) = 180(0.5)t

OpenStudy (anonymous):

Which of the following is an equivalent function to the function below? f(x) = 30(1.03)3x @CBARREDO1

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (anonymous):

HELPPPPPPPP

OpenStudy (anonymous):

The table below shows the functions to represent the sale f(t), in thousands, of four companies in t years: Company A B C D Sale f(t) = 1.05t f(t) = 0.75t f(t) = 1.212t f(t) = 0.9712t Which of the following companies showed an exponential growth in sale?

OpenStudy (anonymous):

forget the other one please do the one above this comment

OpenStudy (anonymous):

@CBARREDO1 @carolinar7

OpenStudy (cbarredo1):

So in order for it to grow, base need to be greater than 1. Make sense? if you repeatedly multiply number, bigger than 1, by itself, it will grow bigger and bigger

OpenStudy (cbarredo1):

Either B or D